Question 3 International Steel Company has budgeted manufacturing overhead costs of $1,995,000. It has allocated overhead on(a) Your answer is correct. Under traditional product costing using machine hours, calculate the total manufacturing cost per(b) Your answer is correct. Under ABC, prepare a schedule showing the calculation of the activity-based overhead rates (per c(c) Calculate the total manufacturing cost per unit for both products under ABC. (Round per cost driver to 2 decimal places.
Answer #1

Calculate manufacturing cost per unit

Soft steel Hard Steel
Direct material 300 200
Direct labor 120 60
Material handling 7*4 = 28 7*6 = 42
Purchase orders 75*2 = 150 225
Product testing 360 480
Machine setup 335 670
Machining 324 324
Total manufacturing cost per unit 1617 2001